OpenStudy (anonymous):

Which of the following is the conjugate of a complex number with 2 as the real part and -8 as the imaginary part? -2 + 8i 2 + 8i 2 - 8i -2 - 8i

OpenStudy (anonymous):

@ganeshie8 2-8i?

ganeshie8 (ganeshie8):

thats your original given complex number not the conjugate

OpenStudy (anonymous):

then would it be 2+8i?

ganeshie8 (ganeshie8):

yes! for conjugate you just flip the sign of imaginary part

OpenStudy (anonymous):

can you check 2 more for me :D?

OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (-3, -4) and (0, 2)? (y - 8) = 2(x - 3) (y + 3) = -2(x + 4) (y - 2) = -2(x + 4) (y + 4) = 2(x + 3)

OpenStudy (anonymous):

i get y=2x+2

ganeshie8 (ganeshie8):

thats right, but they just want the point slope form

ganeshie8 (ganeshie8):

did you get slope = `2` ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

very good, next pick a point and plug them in the point-slope formula

OpenStudy (uri):

y-y1=mx(x-x1) m=2 y1=-4 x1=-3 plug in the formula

ganeshie8 (ganeshie8):

^

OpenStudy (anonymous):

y + 4 = mx(x +3)

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

(y +4) = 2(x+3)

OpenStudy (uri):

Right :D

OpenStudy (anonymous):

ok thank you! and this last one which i forgot how to do Find the average rate of change for the given function from x = -3 to x = 1.

OpenStudy (anonymous):

2, -3, -1, -2 are answer choices

ganeshie8 (ganeshie8):

average rate of change between two points is same as the slope of secant line between those points

ganeshie8 (ganeshie8):

where is your function ?

OpenStudy (anonymous):

can't get picture to post, I'll write it out

OpenStudy (anonymous):

x: -3, 0, 1 y: -3, 3, 5

ganeshie8 (ganeshie8):

simply find the slope between those two given points

ganeshie8 (ganeshie8):

x: ` -3`, 0, `1` y: ` -3`, 3, `5`

ganeshie8 (ganeshie8):

(-3, -3) and (1, 5)

OpenStudy (anonymous):

2

OpenStudy (anonymous):

mx = 2

ganeshie8 (ganeshie8):

Yes! slope of the line joinng both the points is called the average rate of change of function between those two points

ganeshie8 (ganeshie8):

you're done! 2 is the average rate of change

OpenStudy (anonymous):

I didn't know all you had to do was find the slope haha

OpenStudy (anonymous):

thanks again, i just found 2 more that i could use some help with because i forgot the material. do you have time?

ganeshie8 (ganeshie8):

sure, post them as new thread..

OpenStudy (anonymous):

okay